OpenStudy (anonymous):

a regular storage container with an open top is to have a volume of 10 m^3 The length of the base is twice the width material for the base is $10 and the sides $6. find the cost of the materials for the cheapest such container

OpenStudy (anonymous):

im completely lost plz help

OpenStudy (radar):

No wonder, the problem is not clearly stated. For example, is the cost of the base material $10 per sq meter, or is it the total cost? Same goes for the sides, there are 4 sides with 2 of them twice the area of the other two. If it is as stated there is no calculations for the cost,the base $10 plus $6 for the sides, voila total cost $16.00 I don't think that is what they want, but still confusion exists.

OpenStudy (radar):

This may be way out in left field. Let h = height, w=width, 2w=length. Volume=(w)(2w)h=10 cubic meter or h= (10 cubic meter)/2w^2 Now the cost: base =2w^2=$10 sides (there are four) 2 are (wh) each 2 are (2wh) each total sides 2wh+4wh=6$ cost: 2w^2+6wh=$16 now substitute (10 m^3)/(2w^2) for h in the cost equation 2w^2 (6w)(10) ----- + ------ =16 1 2w^2 Simplifying 2w^2 30 ----- + --- =16 1 w 2w^3 +30=16w 2w^3-16w+30=0 dividing thru by 2 w^3-8w=15=0 differentiating 3w^2-8=0 solve and get w= approx 1.8 all other dimensions can now be solved by substitution. I have some doubts about this due to the intermingling units (dollars and meters)